### v1

@@ Abstract
The function of RNA molecules is deeply related to their secondary structure, which determines which nucleobases are accessible for pairing. Most RNA molecules however function through dynamic and heterogeneous structural ensembles. Chemical probing methods (e.g., DMS probing) rely on selective chemical modification of accessible RNA nucleotides to infer base-pairing status, yet the resulting nucleotide-resolution data represent ensemble averages over dynamic RNA conformations. We present MERGE-RNA, a unified, physics-based framework that explicitly models the full experimental pipeline, from the thermodynamics of probe binding to the mutational profiling readout. By integrating measurements across probe concentrations and replicates, our model learns a small set of transferable and interpretable parameters together with minimal sequence-specific soft constraints. This enables the prediction of secondary structure ensembles that best explain the data and the detection of suboptmal structures involved in dynamic processes. We validate MERGE-RNA on diverse RNAs, showing that it achieves strong structural accuracy while preserving essential conformational heterogeneity. In a designed RNA for which we report new DMS data, MERGE-RNA detects transient intermediate states associated with strand displacement, dynamics that remain invisible to traditional methods.
@@
1 Introduction
Figure 1: Schematic overview of the method. Our approach builds a physical model—illustrated on the left side of the figure—that represents the entire pipeline of chemical probing experiments. This includes RNA folding into an ensemble of secondary structures, probe binding, adduct formation, and the final mutational profiling readout. The model is trained on chemical probing data from multiple RNA sequences, probe concentrations, and experimental replicates. From these data, it learns a set of physically meaningful parameters shared across all experiments ($\mu$, $\Delta\mu_{\text{pairing}}$, $p_{\mathrm{bind}}(A)$, $p_{\mathrm{bind}}(C)$, $p_{\mathrm{bind}}(G)$, $p_{\mathrm{bind}}(U)$, $m_{0}$ and $m_{1}$, as defined in the main text). In addition, it estimates sequence-specific soft constraints $\lambda_{i}$ (one per nucleotide position). Once trained, the model can predict the structural ensemble of each RNA at any probe concentration. Importantly, it can also extrapolate to zero probe concentration, providing the unperturbed structural ensemble.
@@ 1 Introduction
RNA molecules play central roles in cellular processes, ranging from catalysis (ribozymes and ribosomes) [doudnaChemicalRepertoireNatural2002, ramakrishnanRibosomeStructureMechanism2002] to gene regulation (including riboswitches and microRNAs) [morrisRiseRegulatoryRNA2014]. This functional versatility arises from their ability to fold into specific secondary and tertiary structures [butcherMolecularInteractionsThat2011]. Importantly, RNA molecules exist as thermodynamic ensembles at both the secondary [boseCausesFunctionsTherapeutic2024] and tertiary [kenRNAConformationalPropensities2023] levels, with function arising from stable conformations, minority states, or dynamic equilibria. While tertiary structure modeling is computationally demanding [buRNAPuzzlesBlindPredictions2025], secondary structure thermodynamic models allow efficient computation of optimal structures [nussinovFastAlgorithmPredicting1980] and base-pairing probabilities [mccaskillEquilibriumPartitionFunction1990]. However, thermodynamic models face significant limitations from fundamental approximations: the nearest-neighbor approximation [tinocoImprovedEstimationSecondary1973], the neglection of pseudoknots [nussinovFastAlgorithmPredicting1980], and the hierarchical folding assumption [tinocoHowRNAFolds1999] may make it difficult to fully capture long-range interactions and context-dependent effects. To address these limitations, researchers have integrated co-evolutionary analysis or chemical probing data. Dimethyl sulfate (DMS) [zaugAnalysisStructureTetrahymena1995, kubotaProgressChallengesChemical2015] has become a powerful tool for structural analysis, providing single-nucleotide resolution data both in vitro and in vivo. DMS reacts preferentially with unpaired adenines and cytosines, with modifications detected as mutations during sequencing (mutational profiling, MaP) [kubotaProgressChallengesChemical2015]. However, interpreting this data is challenging: measured reactivities report on structural ensembles rather than single conformations, and experimental processes introduce systematic biases from reverse transcription and sequencing artifacts. Current approaches typically convert measured reactivities into pseudo-free-energy restraints [deiganAccurateSHAPEdirectedRNA2009], but this can distort predicted ensembles when baseline models already agree with data [washietlRNAFoldingSoft2012], and the coefficients used to compute pseudo-free-energies are based on heuristics rather than physical principles. Despite widespread practice, focusing on a single minimum free-energy (MFE) structure can be misleading as RNA molecules are known to adopt a very heterogeneous ensemble of conformations, of which the MFE structure is often only a minor component. In fact, ignoring structural variability and inter-nucleotide correlations (e.g., contact maps do not fully describe the ensemble) can pose problems both for interpreting RNA function and for properly modeling chemical probing experiments.
@@
In this work we present MERGE-RNA (Multi-system Ensemble Refinement via Generalizable parameters Estimation), a unified physics-based model describes chemical probing experiments with physically meaningful parameters and predicts RNA secondary structure ensembles. Our approach combines a standard thermodynamic folding model with explicit probe-RNA interaction characterized by two key parameters: a chemical potential ($\mu_{r}$) setting the probe concentration scale, and an energetic penalty ($\Delta\mu_{\text{pairing}}$) for probe binding to paired nucleotides. We employ maximum entropy inference to determine sequence-specific soft constraints ($\lambda_{i}$) representing minimal adjustments for model-data consistency, avoiding artificial constraints when baseline predictions already match data. The model explicitly accounts for mutational profiling artifacts, including false positives, false negatives, and position-specific biases. Crucially, our framework enables the simultaneous analysis of data from multiple RNA sequences, probe concentrations, and experimental replicates within a unified model, leveraging the shared physical parameters to improve robustness and extract transferable experimental knowledge.
@@
In the next sections we outline the model and inference procedure, then validate transferability and accuracy on a group of well-characterized structured RNAs. We then apply the framework to the cspA 5 UTR (a known case of temperature-dependent rearrangement). Furthermore, we demonstrate accurate deconvolution of mixed states in synthetic bistable constructs, and finally leverage experimental data to reveal intermediate, suboptimal populations involved in dynamic processes (e.g., strand displacement) that are missed by traditional approaches.
2 Methods
Our approach integrates experimental data from multiple RNA sequences, probe concentrations, and experimental replicates, and outputs the ensemble of secondary structures that are most compatible with the experimental data, as well as the physically meaningful parameters that describe the systems fitted.
2.1 Physical Model
@@
In this section we aim at introducing the physical model underlying our approach, while in the appendix 9.1 we provide a more in depth description and rigorous discussion.
2.1.1 RNA folding and chemical probing
@@
where $z$ is the partition function, $\mathrm{k_{B}}$ the Boltzmann constant and the baseline $F(s)=F_{0}(s)$ is obtained as a sum of thermodynamic parameters. To improve predictions, we refine the ensemble to align with experimental data while minimally perturbing the baseline model. Using the principle of maximum entropy [demartinoIntroductionMaximumEntropy2018], we optimize sequence-specific parameters ($\lambda_{i}$) that act as soft constraints for each nucleotide position [washietlRNAFoldingSoft2012]. These soft constraints modify the baseline ViennaRNA-predicted energy of pairing and remain consistent across experimental replicates and probe concentrations for a given RNA sequence. This results in a structure-dependent energy correction
@@
We separately model the probe reactivity in two steps [calonaciMolecularDynamicsSimulations2023]: a reversible physical binding, described here, followed by an irreversible chemical reaction, as detailed in the next paragraph. At non-zero probe concentrations, we describe physical probe-RNA interactions through a thermodynamic binding model with two trainable parameters ($\mu_{r}$ and $\Delta\mu_{\text{pairing}}$). The reference chemical potential $\mu_{r}$ captures the probability of an unpaired nucleotide to bind to a chemical probe at a reference DMS concentration $\textrm{DMS}]_{r}=1\textrm{M}$. The chemical potential is then corrected based on the actual probe concentration [DMS] as $\mu=\mu_{r}+\mathrm{k_{B}T}\ln\left(\frac{[\text{DMS}]}{[\text{DMS}]_{r}}\right)$. $\Delta\mu_{\text{pairing}}$ differentiates between paired and unpaired sites, so that the effective chemical potential for paired nucleotides becomes $\mu^{\prime}\equiv\mu-\Delta\mu_{\text{pairing}}$. Since chemical probes typically exhibit reduced accessibility to base-paired regions, $\Delta\mu_{\text{pairing}}$ is expected to be positive, though the model could be straightforwardly generalized to cases where this is not true. Incorporating a penalty term for base-paired sites results in a modified partition function. This modification is implemented by applying a concentration-dependent energetic contribution to paired bases $\Delta F_{\text{[DMS]}}(s)$ (derivation is detailed in the appendix, Eq. 14).
@@
Following the physical interaction between a chemical probe and RNA, adducts are formed with probabilities that depend on both nucleobase identity and local RNA structure, under fixed experimental conditions (e.g., temperature and probing time). In the most general framework, this yields eight possible configurations, defined by the combination of nucleotide type (A, U, C, or G) and structural state (paired or unpaired). For DMS probing, due to the position of the potentially reactive nitrogens, adenine and cytosine show negligible reactivity when base-paired, while guanine and uracil retain reduced but structure-independent reactivity. Accordingly, the model is parameterized by four effective binding probabilities, corresponding to the reactive nucleotide-structure combinations: $p_{\text{bind}}(\text{A},\text{unpaired})$, $p_{\text{bind}}(\text{C},\text{unpaired})$, $p_{\text{bind}}(\text{G})$, and $p_{\text{bind}}(\text{U})$.
2.1.2 Mutational Profiling and prediction of reactivity profiles
@@
Following chemical probing, modifications are typically detected through reverse transcription. In the following discussion, we focus on mutational profiling (MaP), though only minor modifications to the model would be required for compatibility with other experimental techniques, such as RT-stop. During reverse transcription, we account for the probability of producing false positive or false negative signals through the parameters $m_{0}$ and $m_{1}$, as detailed in subsubsection 9.1.1. Applying this framework to a given secondary structure $s$, we can compute the expected reactivity profile for that structure, then sum over all secondary structures in the ensemble to obtain the expected reactivity profile for the experiment.
@@
We acknowledge that many sources of systematic error in the experimental pipeline cannot be fully captured by the model. For example, sequencing errors are known to introduce systematic biases in reactivity profiles [liu-weiSequencingAccuracySystematic2024], and reverse transcription can be affected by the presence of secondary structures [verwiltArtifactsBiasesReverse2023]. To account for such effects, we make use of experimental controls (performed without reagent) to estimate position-dependent correction factors $\epsilon_{i}$, which effectively act as a baseline adjustment to the predicted mutation rates at each nucleotide position and are computed so that the prediction at zero concentration of DMS matches the control data. These factors capture systematic biases that are sequence-specific but independent of probe concentration. This allows the model to focus on the concentration-dependent signal in reactivity, which is most informative about RNA structure and probe interactions, while $\epsilon_{i}$ absorbs background effects that remain constant across conditions. In other words, the sequence-dependent $\epsilon_{i}$ parameters model the mutation rate in absence of reagents, whereas the sequence-dependent soft constraints $\lambda_{i}$ affect the pairing population which, in turns, control how much the mutation rate increases when the reagent is added. As discussed in subsubsection 9.1.1, when control experiments are available the parameter $m_{0}$ is omitted, since background correction factors can be directly estimated from the control data. If controls are unavailable, $m_{0}$ is instead fitted to provide a single background correction factor shared across the sequence.
2.2 Parameter Optimization
@@
where $\mathcal{M}_{i}^{\mathrm{exp}}$ is the experimentally observed number of mutations at position $i$. As discussed in the appendix, the overall likelihood of the model across all nucleotide positions is the product of the individual position likelihoods. To facilitate optimization, we minimize the negative log-likelihood, yielding the following loss function:
@@
The derivation and all definitions are expanded in section 9.
This optimization is implemented via a gradient-based algorithm that systematically adjusts model parameters to minimize the discrepancy between predicted and experimental reactivity profiles [byrdLimitedMemoryAlgorithm1995] as implemented in SciPy [virtanenSciPy10Fundamental2020]. The parameter optimization is performed in two consecutive phases, as illustrated in Figure 1. In the first phase, we optimize the global physical parameters that are shared across all sequences and concentrations. These include: the reference chemical potential ($\mu_{r}$), the binding penalty for structured regions ($\Delta\mu_{\text{pairing}}$), the nucleotide-specific adduct formation probabilities ($p_{\text{bind}}(\text{A},\text{unpaired})$, $p_{\text{bind}}(\text{C},\text{unpaired})$, $p_{\text{bind}}(\text{G})$, and $p_{\text{bind}}(\text{U})$), the false negative rate parameter ($m_{1}$), and—when experimental controls are unavailable—the background correction factor ($m_{0}$). This phase optimizes a total of 7 parameters (8 if no controls are available). In the second phase, we look to refine the secondary structure ensemble to find the one that best explains the experimental data. This is achieved by optimizing the sequence-specific soft constrains ($\lambda_{i}$) for each nucleotide position, while keeping the global physical parameters fixed.
@@ 2.3 Experimental procedures
3 Results
3.1 Transferability and accuracy of the physical model

### v2

@@ Abstract
RNA function is deeply tied to secondary structure, with most molecules operating through dynamic and heterogeneous structural ensembles. While current analysis tools typically output single static structures or averaged contact maps, chemical probing methods like DMS capture nucleotide-resolution signals that represent the full structural ensemble, which however remain of difficult structural interpretation. To address this, we present MERGE-RNA, a framework that explicitly describes and outputs RNA as a structural ensemble. By modeling the experimental pipeline through a statistical physics framework, MERGE-RNA learns a small set of transferable and interpretable parameters, enabling the seamless integration of measurements across different molecules, probe concentrations, and replicates in a single optimization to improve robustness. Our model employs a maximum-entropy principle to predict thermodynamic populations, introducing only the minimal sequence-specific adjustments necessary to align the ensemble with experimental data. We validate MERGE-RNA on diverse RNAs, showing that it achieves strong structural accuracy and the ability to fill knowledge gaps in single-conformation reference structures. Furthermore, in a designed RNA construct for which we report new DMS data, MERGE-RNA deconvolves mixed states to reveal transient intermediate populations involved in strand displacement, dynamics that remain invisible to traditional analysis methods.
@@
RNA molecules play central roles in cellular processes, ranging from catalysis (ribozymes and ribosomes) [doudnaChemicalRepertoireNatural2002, ramakrishnanRibosomeStructureMechanism2002] to gene regulation (including riboswitches and microRNAs) [morrisRiseRegulatoryRNA2014]. This functional versatility arises from their ability to fold into specific secondary and tertiary structures [butcherMolecularInteractionsThat2011]. Importantly, RNA molecules exist as thermodynamic ensembles at both the secondary [boseCausesFunctionsTherapeutic2024] and tertiary [kenRNAConformationalPropensities2023] levels, with function arising from stable conformations, minority states, or dynamic equilibria. While tertiary structure modeling is computationally demanding [buRNAPuzzlesBlindPredictions2025], secondary structure thermodynamic models allow efficient computation of optimal structures [nussinovFastAlgorithmPredicting1980] and base-pairing probabilities [mccaskillEquilibriumPartitionFunction1990]. However, thermodynamic models face significant limitations from fundamental approximations: the nearest-neighbor approximation [tinocoImprovedEstimationSecondary1973], the neglect of pseudoknots [nussinovFastAlgorithmPredicting1980], and the hierarchical folding assumption [tinocoHowRNAFolds1999] may make it difficult to fully capture long-range interactions and context-dependent effects. To address these limitations, researchers have integrated co-evolutionary analysis or chemical probing data. Dimethyl sulfate (DMS) [zaugAnalysisStructureTetrahymena1995, kubotaProgressChallengesChemical2015] has become a powerful tool for structural analysis, providing single-nucleotide resolution data both in vitro and in vivo. DMS reacts preferentially with unpaired adenines and cytosines, with modifications detected as mutations during sequencing (mutational profiling, MaP) [kubotaProgressChallengesChemical2015]. However, interpreting this data is challenging: measured reactivities report on structural ensembles rather than single conformations, and experimental processes introduce systematic biases from reverse transcription and sequencing artifacts. Current approaches typically convert measured reactivities into pseudo-free-energy restraints [deiganAccurateSHAPEdirectedRNA2009], but this modifies predicted ensembles even when baseline models already agree with data [washietlRNAFoldingSoft2012], and the coefficients used to compute pseudo-free-energies are based on heuristics rather than physical principles.
Figure 1: Schematic overview of the method. Our approach builds a physical model—illustrated on the left side of the figure—that represents the entire pipeline of chemical probing experiments. This includes RNA folding into an ensemble of secondary structures, probe binding, adduct formation, and the final mutational profiling readout. The model is trained on chemical probing data from multiple RNA sequences, probe concentrations, and experimental replicates. From these data, it learns a set of physically meaningful parameters shared across all experiments ($\mu_{r}$, $\Delta\mu_{\text{pairing}}$, $p_{\mathrm{bind}}(A)$, $p_{\mathrm{bind}}(C)$, $p_{\mathrm{bind}}(G)$, $p_{\mathrm{bind}}(U)$, $m_{0}$ and $m_{1}$, as defined in the main text). In addition, it estimates sequence-specific soft constraints $\lambda_{i}$ (one per nucleotide position). Once trained, the model can predict the structural ensemble of each RNA at any probe concentration. Importantly, it can also extrapolate to zero probe concentration, providing the unperturbed structural ensemble.
Despite widespread practice, focusing on a single minimum free-energy (MFE) structure can be misleading as RNA molecules are known to adopt an ensemble of conformations, of which the MFE structure is often only a minor component. In fact, ignoring structural heterogeneity can pose problems both for interpreting RNA function and for properly modeling chemical probing experiments.
@@
In this work we present MERGE-RNA (Multi-system Ensemble Refinement via Generalizable parameters Estimation), a unified physics-based model that describes chemical probing experiments with physically meaningful parameters and predicts RNA secondary structure ensembles. Our approach combines a standard thermodynamic folding model with explicit probe-RNA interactions characterized by two key parameters: a chemical potential ($\mu_{r}$) setting the probe concentration scale, and an energetic penalty ($\Delta\mu_{\text{pairing}}$) for probe binding to paired nucleotides. We employ maximum entropy inference to determine sequence-specific soft constraints ($\lambda_{i}$) representing minimal adjustments for model-data consistency, avoiding artificial constraints when baseline predictions already match data. The model explicitly accounts for mutational profiling artifacts, including false positives, false negatives, and position-specific biases. Crucially, our framework enables the simultaneous analysis of data from multiple RNA sequences, probe concentrations, and experimental replicates within a unified model, leveraging the shared physical parameters to improve robustness and extract transferable experimental knowledge.
@@
In the next sections we outline the model and inference procedure, then validate transferability and accuracy on a group of well-characterized structured RNAs. We then apply the framework to the cspA 5 untranslated region (UTR) (a known case of temperature-dependent rearrangement). Furthermore, we demonstrate accurate deconvolution of mixed states in synthetic bistable constructs, and finally leverage experimental data to reveal intermediate, suboptimal populations involved in dynamic processes (e.g., strand displacement) that are missed by traditional approaches.
Methods
Physical Model
@@ Methods
In this section we aim at introducing the physical model underlying our approach, while in the SI Appendix we provide a more in-depth description and rigorous discussion.
RNA folding and chemical probing
@@
where $z$ is the partition function, $\mathrm{k_{B}}$ is the Boltzmann constant, and the baseline $F(s)=F_{0}(s)$ is obtained as a sum of thermodynamic parameters. To improve predictions, we refine the ensemble to align with experimental data while minimally perturbing the baseline model. Using the principle of maximum entropy [demartinoIntroductionMaximumEntropy2018], we optimize sequence-specific parameters ($\lambda_{i}$) that act as soft constraints for each nucleotide position [washietlRNAFoldingSoft2012]. These soft constraints modify the baseline ViennaRNA-predicted energy of pairing and remain consistent across experimental replicates and probe concentrations for a given RNA sequence. This results in a structure-dependent energy correction
@@
We separately model the probe reactivity in two steps [calonaciMolecularDynamicsSimulations2023]: a reversible physical binding, described here, followed by an irreversible chemical reaction, as detailed in the next paragraph. At non-zero probe concentrations, we describe physical probe-RNA interactions through a thermodynamic binding model with two trainable parameters ($\mu_{r}$ and $\Delta\mu_{\text{pairing}}$). The reference chemical potential $\mu_{r}$ captures the probability of an unpaired nucleotide to bind to a chemical probe at a reference DMS concentration $\textrm{DMS}]_{r}=1\textrm{M}$. The chemical potential is then corrected based on the actual probe concentration [DMS] as $\mu=\mu_{r}+\mathrm{k_{B}T}\ln\left(\frac{[\text{DMS}]}{[\text{DMS}]_{r}}\right)$. $\Delta\mu_{\text{pairing}}$ differentiates between paired and unpaired sites, so that the effective chemical potential for paired nucleotides becomes $\mu^{\prime}\equiv\mu-\Delta\mu_{\text{pairing}}$. Since chemical probes typically exhibit reduced accessibility to base-paired regions, $\Delta\mu_{\text{pairing}}$ is expected to be positive, though the model could be straightforwardly generalized to cases where this is not true. Incorporating a penalty term for base-paired sites results in a modified partition function. This modification is implemented by applying a concentration-dependent energetic contribution to paired bases $\Delta F_{\text{[DMS]}}(s)$ (derivation is detailed in the SI Appendix).
@@
Following the physical interaction between a chemical probe and RNA, adducts are formed with probabilities that depend on both nucleobase identity and local RNA structure, under fixed experimental conditions (e.g., temperature and probing time). In the most general framework, this yields eight possible configurations, defined by the combination of nucleotide type (A, U, C, or G) and structural state (paired or unpaired). For DMS probing, due to the position of the potentially reactive nitrogens, adenine and cytosine show negligible reactivity when base-paired, while guanine and uracil retain reduced but secondary structure-independent reactivity. Accordingly, the model is parameterized by four effective binding probabilities, corresponding to the reactive nucleotide-structure combinations: $p_{\text{bind}}(\text{A},\text{unpaired})$, $p_{\text{bind}}(\text{C},\text{unpaired})$, $p_{\text{bind}}(\text{G})$, and $p_{\text{bind}}(\text{U})$.
Mutational Profiling and prediction of reactivity profiles
@@
Following chemical probing, modifications are typically detected through reverse transcription. Here, we focus on mutational profiling (MaP), though only minor modifications to the model would be required for compatibility with other experimental techniques, such as RT-stop. During reverse transcription, we account for the probability of producing false positive or false negative signals through the parameters $m_{0}$ and $m_{1}$, as detailed in the SI Appendix. Applying this framework to a given secondary structure $s$, we can compute the expected reactivity profile for that structure, then sum over all secondary structures in the ensemble to obtain the expected reactivity profile for the experiment.
@@
Sequencing errors are known to introduce systematic biases in reactivity profiles [liu-weiSequencingAccuracySystematic2024], and reverse transcription can be affected by the presence of secondary structures [verwiltArtifactsBiasesReverse2023]. To account for such effects, we make use of experimental controls (performed without reagent) to estimate position-dependent correction factors $\epsilon_{i}$, which effectively act as a baseline adjustment to the predicted mutation rates at each nucleotide position and are computed so that the prediction at zero concentration of DMS matches the control data. These factors capture systematic biases that are sequence-specific but independent of probe concentration. This allows the model to focus on the concentration-dependent signal in reactivity, which is most informative about RNA structure and probe interactions, while $\epsilon_{i}$ absorbs background effects that remain constant across conditions. In other words, the sequence-dependent $\epsilon_{i}$ parameters model the mutation rate in absence of reagents, whereas the sequence-dependent soft constraints $\lambda_{i}$ affect the pairing population which, in turn, controls how much the mutation rate increases when the reagent is added. As discussed in the SI Appendix, when control experiments are available the parameter $m_{0}$ is omitted, since background correction factors can be directly estimated from the control data. If controls are unavailable, $m_{0}$ is instead fitted to provide a single background correction factor shared across the sequence.
Parameter Optimization
@@
where $\mathcal{M}_{i}^{\mathrm{exp}}$ is the experimentally observed number of mutations at position $i$. As discussed in the SI Appendix, the overall likelihood of the model across all nucleotide positions is the product of the individual position likelihoods. To facilitate optimization, we minimize the negative log-likelihood, yielding the following loss function:
@@
The derivation and all definitions are expanded in the SI Appendix.
This optimization is implemented via a gradient-based algorithm that systematically adjusts model parameters to minimize the discrepancy between predicted and experimental reactivity profiles [byrdLimitedMemoryAlgorithm1995] as implemented in SciPy [virtanenSciPy10Fundamental2020]. The parameter optimization is performed in two consecutive phases, as illustrated in Figure 1. In the first phase, we optimize the global physical parameters that are shared across all sequences and concentrations. These include: the reference chemical potential ($\mu_{r}$), the binding penalty for structured regions ($\Delta\mu_{\text{pairing}}$), the nucleotide-specific adduct formation probabilities ($p_{\text{bind}}(\text{A},\text{unpaired})$, $p_{\text{bind}}(\text{C},\text{unpaired})$, $p_{\text{bind}}(\text{G})$, and $p_{\text{bind}}(\text{U})$), the false negative rate parameter ($m_{1}$), and—when experimental controls are unavailable—the background correction factor ($m_{0}$). This phase optimizes a total of 7 parameters (8 if no controls are available). In the second phase, we look to refine the secondary structure ensemble to find the one that best explains the experimental data. This is achieved by optimizing the sequence-specific soft constraints ($\lambda_{i}$) for each nucleotide position, while keeping the global physical parameters fixed.
@@
Results
Transferability and accuracy of the physical model